OpenStudy (anonymous):

For the following quadratic, find the vertex, the vertical intercept, and the horizontal intercepts y(x) = 2x^2+10x+12

OpenStudy (anonymous):

I know that I can factor out the equation to get my vertex, but I can only factor as far as a 2 so I have 2(x^2+5x+6)

OpenStudy (anonymous):

How do I factor that out more?

OpenStudy (anonymous):

1. f(x) = x^2 - 6x + 8 y = f(x) so write it: y = x^2 - 6x + 8

OpenStudy (anonymous):

Oh, it should be 2x^2 so that I can factor out a 2

OpenStudy (anonymous):

2x^2+10x+12 is the equation

OpenStudy (anonymous):

Actually you can type the question in on algebra.com here's the link :) http://www.algebra.com/algebra/homework/Functions/Functions.faq.question.145531.html

OpenStudy (anonymous):

With explanation let me know if that helps ..

OpenStudy (anonymous):

I'll give that a try. Thanks :)

OpenStudy (anonymous):

I asked because I want to get better at factoring.

OpenStudy (anonymous):

well it gives a good explanation let me know if it helps

OpenStudy (anonymous):

@whpalmer4 he is better at explaining this than me

OpenStudy (anonymous):

That did help me out Shay

OpenStudy (anonymous):

I like the first explanation so I don't have to remember formulas :)

OpenStudy (anonymous):

Ok great your welcome :)